How do I show that (cscO + cotO)(cscO - cotO) = 1 Note: All of the O have a horizontal line going through the middle. They are NOT zeros. PLEASE HELP ME! And If you could can please go step by step because that is the only way I will learn. Thank you :)
Answer by stanbon(75887)

(cscO + cotO)(cscO - cotO) = 1
----
You have the form: (a+b)(a-b) = a^2-b^2
-------
(csc^2 - cot^2) = 1
-----
((1/sin^2) - (cos^2/sin^2) = 1
----
(1-cos^2)/sin^2 = 1
-----
sin^2/sin^2 = 1
----
1 = 1
=============
Cheers,
Stan H.